OpenStudy (anonymous):

What is the value of the expression –c + d when c = –3.5 and d = –12.8? A. –9.3 B. –16.3 C. 15.3 D. 16.3

OpenStudy (anonymous):

B

OpenStudy (anonymous):

+3.5-12.8 = -9.3

OpenStudy (anonymous):

@Learner11 incorrect

OpenStudy (anonymous):

−3.5+−12.8= -16.3

OpenStudy (anonymous):

-c+d =- (-3.5)+(-12.8) =+3.5-12.8 = -9.3

OpenStudy (anonymous):

@BlooooooBlahBLIIIIK ?

OpenStudy (anonymous):

substitute -3.5 for c and -12.8 for d then do the math thats all you have to do when they give substitution numbers in the question.

OpenStudy (anonymous):

am I right guys ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks . You can give me a medal :)

OpenStudy (anonymous):

lol yeah :)

OpenStudy (anonymous):

Learner11 :P

OpenStudy (anonymous):

whats ur name

OpenStudy (anonymous):

I told :P

OpenStudy (anonymous):

cool u play 20 question

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

do u play the game 20 questions

OpenStudy (anonymous):

nope .

OpenStudy (anonymous):

actualy it is 16.3 i redid my math it is b

OpenStudy (anonymous):

But How ?

OpenStudy (anonymous):

look if u substitute -3.5 for c and - 12.8 for d that would give u the equation -3.5+(-12.8 which =16.3 add it up in a calculator and you willl see